To solve systems of equations by the substitution method. Please help my son to solve these 3 homework questions:
(1) x – y = 2
4x – 3y = 11
(2) – 2x + 3y = 14
x + 2y = 7
(3) 6x – y = -4
2x + 2y = 15
Thanks so much for all your help.

Answer by stanbon(75887)

(1) x – y = 2
4x – 3y = 11
Solve the 1st equation for x: x=y+2
Substitue in the 2nd equation to solve for y:
4(y+2)-3y=11
4y+8-3y=11
y=3
--------
Substitue into x=y+2 to solve for x:
x=3+2
x=5
--------
SOLUTION:
x=5, y=3
======================
(2) – 2x + 3y = 14
x + 2y = 7
Solve the 2nd equation for x:
x=-2y+7
Substitute into the 1st equation:
-2(-2y+7)+3y=14
4y-14+3y=14
7y=28
y=4
-------
Substitute into x=-2y+7 to solve for x:
x=-2(4)+7
x=-1
---------
SOLUTION:
x=-1, y=4
========================
(3) 6x – y = -4
2x + 2y = 15
solve the 1st for y to get:
y=6x+4
Substitute into the 2nd to get:
2x+2(6x+4)=15
2x+12x+8=15
14x=7
x=1/2
-----
Substitute into y=6x+4 to get:
y=6(1/2)+4
y=3+4
y=7
---------
Solution:
x=1/2; y=7
Cheers,
Stan H.